# Assimilative Causal Inference

### v1

@@ D.3 A Stochastic Model Capturing the El Niño-Southern Oscillation (ENSO) Diversity
Figure 4.5: ACI values and CIRs for some model-simulated EP events (EEP EN and MEP EN stand for extreme and moderate EP El Niño, respectively). Panel (a): Reconstructed spatiotemporal pattern of SST from the time series $T_{C}$ and $T_{E}$. The wind profile, plotted above the dateline, is superimposed, along with the decadal variable $I$, which is also centered at the dateline. Positive $\tau$ (indicating westerly winds) appear to the right of the dateline, while negative values appear to the left. A weakening Walker circulation appears close to the dateline, while a stronger $I$ appears further to its right. Panel (b): ACI values from $T_{C}$ to $T_{E}$. Panel (c): Time series of $T_{C}$ and the corresponding CIR (depicted as whiskers) from $T_{C}$ to $T_{E}$. Panels (d)-–(e) and (f)-–(g): Similar to (a)-–(b) but for $h_{W}$ and $\tau$. The two rows represent two distinct periods. The Walker circulation in both cases is nearly zero, which is the usual condition to trigger EP events.
@@
Figure 4.6 shows ACI values and CIRs for model-simulated CP events (target variable: $T_{C}$). When the decadal variable $I$ is significant, $u$ contributes substantially to CP events through the enhanced zonal advective feedback ($\sigma(I)\propto I$) according to (4.3c), peaking slightly before CP El Niño events. The WP thermocline ($h_{W}$) also affects $T_{C}$ via positive feedback in (4.3c), but peaks earlier than the CP event. While the $u$ and $h_{W}$ ACIs sometimes coincide due to state-space correlations, $h_{W}$’s CIR persists longer due to WP-CP information transfer delays. Wind ($\tau$) influences CP events on shorter intraseasonal timescales. Unlike EP events, CP events show balanced contributions from $u$, $h_{W}$, and $\tau$. The strong ACI value from $u$ to $T_{C}$ around $t=1985$ corresponds to the rapid transition from CP El Niño to La Niña, where the value of $u$ changes dramatically from positive to negative. Furthermore, during the second year of the multi-year CP La Niña events (i.e., $t=1986$), all three ACIs drop suddenly, as the signals of all these three variables remain near zero. This corresponds to a typical discharge phase of ENSO.
@@
Figure 4.6: ACI values and CIRs for some model-simulated CP events (EEP EN and MEP EN stand for extreme and moderate EP El Niño, respectively). Panel (a): Reconstructed spatiotemporal pattern of SST from the time series $T_{C}$ and $T_{E}$. The wind profile, plotted above the dateline, is superimposed, along with the decadal variable $I$, which is also centered at the dateline. Positive $\tau$ (indicating westerly winds) appear to the right of the dateline, while negative values appear to the left. A weakening Walker circulation appears close to the dateline, while a stronger $I$ appears further to its right. Panel (b): ACI values from $u$ to $T_{C}$. Panel (c): Time series of $T_{C}$ and the corresponding CIR (depicted as whiskers) from $T_{C}$ to $T_{E}$. Panels (d)-–(e) and (f)-–(g): Similar to (a)-–(b) but for $h_{W}$ and $\tau$. The Walker circulation in this case is strong, which is the usual condition to trigger CP events.
Supporting Information References
  • [1] H. Abou-Kandil, G. Freiling, V. Ionescu, and G. Jank (2003) Matrix Riccati Equations in Control and Systems Theory. Birkhäuser Basel. External Links: ISBN 9783034880817, ISSN 2324-9757, Link, Document Cited by: §C.2.
  • [2] M. Andreou, N. Chen, and Y. Li (2024) An adaptive online smoother with closed-form solutions and information-theoretic lag selection for conditional gaussian nonlinear systems. arXiv preprint arXiv:2411.05870. Cited by: §A.5, §B.2, Theorem B.3.
  • [3] D. Cai, R. Kleeman, and A. Majda (2002) A Mathematical Framework for Quantifying Predictability Through Relative Entropy. Methods and Applications of Analysis 9 (3), pp. 425–444. External Links: ISSN 1945-0001, Link, Document Cited by: §A.3.
  • [4] N. Chen, X. Fang, and J. Yu (2022) A multiscale model for El Niño complexity. npj Climate and Atmospheric Science 5 (1), pp. 16. Cited by: §D.3, §D.3.
  • [5] N. Chen and A. J. Majda (2018) Conditional Gaussian systems for multiscale nonlinear stochastic systems: Prediction, state estimation and uncertainty quantification. Entropy 20 (7), pp. 509. Cited by: §B.1.1.
  • [6] D. Cooke and R. W. Hiorns (1981) The mathematical theory of the dynamics of biological populations 2: based on the proceedings of a conference on the mathematical theory of the dynamics of biological populations organised by the institute of mathematics and its applications and held in oxford, 1st-3rd july, 1980. Academic Press. Cited by: §D.2.
  • [7] J. H. Gallier (2019) The Schur Complement and Symmetric Positive Semidefinite (and Definite) Matrices. Note: https://www.cis.upenn.edu/~jean/schur-comp.pdf[Online; Accessed 25-March-2025] Cited by: §C.2, §C.2.
  • [8] X. Han and P. E. Kloeden (2017) Random Ordinary Differential Equations and Their Numerical Solution. Springer Singapore. External Links: ISBN 9789811062650, ISSN 2199-3149, Link, Document Cited by: §C.2.
  • [9] N. Ikeda and S. Watanabe (2014) Stochastic differential equations and diffusion processes. Vol. 24, Elsevier. Cited by: §A.2.1.
  • [10] F. Jin (1997) An equatorial ocean recharge paradigm for ENSO. Part I: Conceptual model. Journal of the atmospheric sciences 54 (7), pp. 811–829. Cited by: §D.3.
  • [11] R. E. Kalman and R. S. Bucy (1961-03) New Results in Linear Filtering and Prediction Theory. Journal of Basic Engineering 83 (1), pp. 95–108. External Links: ISSN 0021-9223, Link, Document Cited by: §B.1.1.
  • [12] R. Kleeman (2002) Measuring Dynamical Prediction Utility Using Relative Entropy. Journal of the Atmospheric Sciences 59 (13), pp. 2057 – 2072. External Links: Document, Link Cited by: §A.3.
  • [13] R. Kleeman (2011-03) Information Theory and Dynamical System Predictability. Entropy 13 (3), pp. 612–649. External Links: ISSN 1099-4300, Link, Document Cited by: §A.3.
  • [14] X. S. Liang (2016) Information flow and causality as rigorous notions ab initio. Physical Review E 94 (5), pp. 052201. Cited by: SI. C.
  • [15] R. S. Liptser and A. N. Shiryaev (2001) Statistics of random processes I: General theory & Statistics of Random Processes II: Applications. Springer Science & Business Media. Cited by: §A.4.1, §B.1.1, §B.1, Theorem B.1, Theorem B.2, §C.1.
  • [16] A. J. Majda and J. Harlim (2012) Physics constrained nonlinear regression models for time series. Nonlinearity 26 (1), pp. 201. Cited by: §D.1.
  • [17] A. J. Majda, I. Timofeyev, and E. Vanden-Eijnden (2003) Systematic strategies for stochastic mode reduction in climate. Journal of the Atmospheric Sciences 60 (14), pp. 1705–1722. Cited by: §B.1.1.
  • [18] T. Neckel and F. Rupp (2013-11) Random Differential Equations in Scientific Computing. Versita. External Links: ISBN 9788376560243, Link, Document Cited by: §C.2.
  • [19] B. Oksendal (2003-07) Stochastic Differential Equations. Universitext, Springer, Berlin, Germany (en). Cited by: §A.2.1.
  • [20] H. E. Rauch, F. Tung, and C. T. Striebel (1965) Maximum likelihood estimates of linear dynamic systems. AIAA journal 3 (8), pp. 1445–1450. Cited by: §B.1.1.
  • [21] B. L. Rozovsky and S. V. Lototsky (2018) Stochastic Evolution Systems: Linear Theory and Applications to Non-Linear Filtering. Vol. 89, Springer Science & Business Media. Cited by: §A.2.1, §A.2.1, §A.2.1, §A.2.3, §A.4.1.
  • [22] J. Vialard, F. Jin, M. McPhaden, A. Fedorov, W. Cai, S. An, D. Dommenget, X. Fang, M. Stuecker, C. Wang, et al. (2025) The El Niño Southern Oscillation (ENSO) recharge oscillator conceptual model: Achievements and future prospects. Reviews of Geophysics 63 (1), pp. e2024RG000843. Cited by: §D.3.

### v2

@@ D.3.2 Additional Results Using Model-Generated Data
Supplementary Figure 5: Conditional ACI values and CCIRs for some model-simulated EP events (EEP EN and MEP EN stand for extreme and moderate EP El Niño, respectively, and LN for La Niña); $t\in[364,368]$ and $t\in[845,849]$. Panel (a): Hovmöller diagram of the SST anomalies in the equatorial Pacific reconstructed from $T_{C}$ and $T_{E}$ via a spatiotemporal multivariate regression technique [chen2022multiscale]. The wind profile $\tau$, plotted above and centered around the dateline, is superimposed along with the decadal variable II, whose time series is also positioned relative to the dateline: Positive $\tau$ (indicating westerly winds) appear to the right of the dateline, while negative values appear to the left, and a weakening Walker circulation appears close to the dateline, while a stronger II appears further to its right. Panel (b): Conditional ACI values from $T_{C}$ to $T_{E}$. Panel (c): Time series of $T_{C}$ and the corresponding CCIRs from $T_{C}$ to $T_{E}$ (depicted as whiskers at each month, similar to Panel (a) of Supplementary Figure 2). Panels (d)–(e) and (f)–(g): Similar to (a)–(b) but for $h_{W}$ and $\tau$. The two rows represent two distinct periods. The Walker circulation in both cases is nearly zero, which is the usual condition to trigger EP events.
@@
Supplementary Figure 6 shows conditional ACI values and CCIRs for model-simulated CP events (target variable: $T_{C}$). When the decadal variable $I$ is significant, $u$ contributes substantially to CP events through the enhanced zonal advective feedback ($\sigma(I)\propto I$) according to (4.3c), peaking slightly before CP El Niño events. The WP thermocline ($h_{W}$) also affects $T_{C}$ via positive feedback in (4.3c), but peaks earlier than the CP event. While the $u$ and $h_{W}$ ACIs sometimes coincide due to state-space correlations, $h_{W}$’s CIR persists longer due to WP-CP information transfer delays. Wind ($\tau$) influences CP events on shorter intraseasonal timescales. Unlike EP events, CP events show balanced contributions from $u$, $h_{W}$, and $\tau$. The strong ACI value from $u$ to $T_{C}$ around $t=1985$ corresponds to the rapid transition from CP El Niño to La Niña, where the value of $u$ changes dramatically from positive to negative. Furthermore, during the second year of the multi-year CP La Niña events (i.e., $t=1986$), all three ACIs drop suddenly, as the signals of all these three variables remain near zero. This corresponds to a typical discharge phase of ENSO.
@@
Supplementary Figure 6: Conditional ACI values and CCIRs for some model-simulated CP events (CP EN and MEP EN stand for CP and moderate EP El Niño, respectively, and LN for La Niña); $t\in[1983.5,1987.5]$ and $t\in[2087,2091]$. Panel (a): Same as Panel (a) of Figure 5. Panel (b): Conditional ACI values from $u$ to $T_{C}$. Panel (c): Time series of $u$ and the corresponding CCIRs from $u$ to $T_{C}$ (depicted as whiskers at each month, similar to Panel (a) of Supplementary Figure 2). Panels (d)–(e) and (f)–(g): Similar to (a)–(b) but for $h_{W}$ and $\tau$. The Walker circulation in this case is strong, which is the usual condition to trigger CP events.
D.3.3 ACI on Real-World ENSO Data
In this subsection, we apply the ACI framework on real-world observational data to study the ENSO diversity through (4.3).
The oceanic data used here, including the monthly SST, zonal current velocity, and thermocline depth (calculated from the potential temperature at the depth of the 20C isotherm) measurements, are sourced from the GODAS dataset (https://www.psl.noaa.gov/data/gridded/data.godas.html) [behringer2004evaluation]. The present study solely uses oceanic anomalies, calculated by removing the monthly mean climatology over the whole analysis period. The data spans from January 1982 to December 2017, chosen to fall within the satellite era when observations are more reliable. The atmospheric wind data is obtained from the National Centers for Environmental Prediction-National Center for Atmospheric Research (NCEP-NCAR) reanalysis data set (https://psl.noaa.gov/data/gridded/data.ncep.reanalysis.html) [kalnay2018ncep] over the same period.
The model used within ACI remains the stochastic conceptual model defined in (4.3a)–(4.3f). However, because the observations are now drawn from the real world rather than from synthetic data, this study no longer has a perfect-model setup. Nevertheless, as shown in [chen2022multiscale], the conceptual model in (4.3a)–(4.3f) successfully reproduces key statistical features of observed ENSO variability and generates time series that are qualitatively consistent with real ENSO episodes. Hence, the model error is expected to be relatively small.
Daily data are available for the wind burst amplitudes $\tau$, whereas all other variables are monthly. To ensure the stability of the filter, smoother, and online-smoother algorithms required for the conditional ACI analysis and the computation of the CCIRs, we linearly interpolate all data onto a uniform temporal grid with a spacing of approximately 7 hours and 20 minutes. Linear interpolation has been shown to be more compatible with data assimilation than smoother interpolation methods [harlim2011interpolating]. Since the oceanic variables primarily evolve on interannual timescales, this approximation is not expected to alter the data’s underlying dynamics. This procedure also provides an opportunity to test the robustness of the ACI framework with respect to potential numerical discretization errors.
Supplementary Figure 7: Conditional ACI values and CCIRs for some real-world EP events; Similar to Supplementary Figure 5 but the observations are collected from real-world data.
Supplementary Figure 8: Conditional ACI values and CCIRs for some real-world CP events; Similar to Supplementary Figure 6 but the observations are collected from real-world data.
Supplementary Figure 7 includes the historically strongest EP El Niño event in 1998 and a moderately strong EP event in 1987, together with their subsequent La Niña phases. Supplementary Figure 8 presents several CP events that occurred in the early 21st century. The results based on real-world observations show qualitatively similar conclusions to those in Supplementary Figures 56, which use model-generated synthetic data.
Specifically, during EP El Niño phases, the causal links $T_{C}\rightarrow T_{E}$, $h_{W}\rightarrow T_{E}$, and $\tau\rightarrow T_{E}$ are identified. In addition, the thermocline continues to play a key role during EP La Niña events through the recharge–discharge mechanism, while there are no clear causal influences from $T_{C}$ and $\tau$ to $T_{E}$, consistent with the classical discharge process. Likewise, during CP El Niño phases, $u\rightarrow T_{C}$ and $\tau\rightarrow T_{C}$ emerge as the two most significant causal relationships, while during CP La Niña phases $h_{W}\rightarrow T_{C}$ is also detected. All of these findings are consistent with those obtained from the model-generated experiments.
The results presented here suggest that ACI performs reasonably well in the presence of noisy observations, moderate model error, and some numerical discretization effects. While more systematic studies are needed to fully quantify the influence of these factors, the results here provide some preliminary evidence that the ACI framework maintains a degree of robustness under these conditions.
Finally, it is worth emphasizing the distinction between ACI and many baseline causal inference methods for analyzing real-world ENSO datasets [risser2025granger, runge2023causal, falasca2024data]. Most of the current approaches are purely data-driven and focus on identifying broad, time-averaged cause-and-effect relationships. Their findings are generally consistent with physical mechanisms, such as the causal influence of the thermocline depth on the SST and the discharge–recharge loop linking El Niño and La Niña. Yet, ACI provides a more refined and event-specific indicator that reveals how individual variables contribute to the causal structure of each ENSO episode, an aspect often missing from state-of-the-art techniques. ACI yields markedly different causal strengths and CIR patterns across events, providing valuable insight into instantaneous, time-localized cause-and-effect behavior. As a natural direction for future work, the ACI framework can be coupled with a range of ENSO models and expanded to include additional variables, thereby enhancing causal discovery with a particular focus on instantaneous relationships.
Supplementary Information References